# A Learnable Prior Improves Inverse Tumor Growth Modeling

### v1

@@ 3.2.2 Estimated Parameter Comparison
Similar results to DL and DL ensemble are found for subsequent evolutionary sampling (Table 1). Limited information, introduced by stochastic thresholding of test data, might lead to a performance threshold that is impossible to overcome theoretically. For the combined version of sampling and prior, a slightly worse result is found for $\mu_{D}$ and $\mu_{\rho}$, whereas the origin deviation is comparable to the ensemble mean.
Even though parameter estimation does not improve with the network ensemble and subsequent evolutionary sampling, the rationale for training the ensemble is founded on the anticipation of achieving more stable predictions and providing variance estimation for the DL-Prior to optimally initialize the subsequent sampling. This premise is examined by comparing the resulting tumor concentrations in the following section.
@@ 3.2.3 Estimated Tumor Concentration
Subsequent sampling decreases the error to the ground truth of the estimated tumor concentration as shown in Table 2. This highlights the synergistic combination of sampling with a DL-Prior. Further, the combined version of CMA-ES sampling and DL-Prior archives a significantly better Dice-score than the DL-Prior alone over the whole range of tumor concentrations (Figure 5). Within the clinically relevant range, up to 40% tumor concentration, a Dice overlap of 95% is found, further demonstrating the superiority of combining DL with traditional sampling.
@@ 3.2.4 Runtime comparison
To guarantee the integrity of our comparative analysis regarding time, we have included performance metrics for the specific subset of samples utilized in the temporal comparison, as detailed in Table 3. Sampling enhanced with a DL-Prior demonstrates improved results in terms of the Dice. However, the observed differences between these methods fall within the range of their respective error margins. These observations suggest that both approaches converge towards a similar minimum, legitimizing the time comparison of Figure 6.
Figure 6: The mean convergence time of our DL-inference (orange) and our combination of classical sampling with DL-Prior (blue) is compared to sampling without prior (purple, dashed).
Table 3: Tumor concentration comparison between DL-Prior with subsequent sampling and sampling without a prior for a subset of patients. The Dice-score at different thresholds (0.01, 0.1, 0.5) is given. Simulation time differences are shown in Figure 6.
@@ 4 Validation on Patient Data
The quantitative ability of different methods to fit real tumor segmentations is compared in Table 4. The Dice-score is compared for CET and edema regions. It is clearly visible that sampling improves the simulation fit dramatically. The differences in Dice-score between naive and prior included sampling are minor, while a small advantage is on the side of naive sampling. Regarding the runtime, a clear advantage is seen for DL-Piror + sampling.
Figure 7: Real patient example with different inversion strategies. The first image shows the ground truth tumor segmentation before surgery. Light green shows the edema region, while dark green represents the CET, which is included in the following images as a black outline. The mean likelihood development over time comparing DL-Prior sampling to naive sampling is plotted. The margin shows the minimum and maximum values for each CMA-ES epoch, and the vertical lines depict the time points of convergence.
Table 4: For nine real patients, the Dice-score to the tumor segmentation is compared for different algorithms, thresholded at the $\overline{c^{th}}=\{0.25,0.675\}$. Additionally, the convergence time is reported for each method.
@@
Naive Sampling 0.70 $\pm$ 0.03 0.70 $\pm$ 0.02 (16.3 $\pm$ 1.7) h
DL 0.39 $\pm$ 0.06 0.44 $\pm$ 0.04 $<$ 1 min
DL Ensemble 0.44 $\pm$ 0.08 0.47 $\pm$ 0.05 $<$ 1 min
LMI 0.34 $\pm$ 0.09 0.55 $\pm$ 0.04 $<$ 1 min
@@ 5 Discussion
5.1 Deep Learning
A new DL-based method is introduced, operating solely in patient image space compared to previous atlas-based methods [13] by explicitly incorporating patient anatomy into the network. With our method, a significantly better parameter estimation is found (Table 1). Additionally, a better Dice overlap to the ground truth concentration is observed (Figure 5). We assume that the abundance of registration errors between the patient and the atlas space causes this improvement.
@@ 6 Outlook
Finally, an evaluation of large-scale clinical data is needed. As the availability of longitudinal data is highly limited due to the rapid surgery after detection for most patients, other validation metrics are considered. Prediction of tumor recurrence represents one of the most promising approaches. Besides the influence of treatment procedures like surgery and radiotherapy, it was shown that a certain degree of recurrence prediction is possible [37, 38]. Showing the ability to predict recurrence with a physically constrained model would provide an excellent argument to proceed with clinical trials testing model-based radiotherapy planning.

### v2

@@ 3.2.2 Estimated Parameter Comparison
Similar results to DL and DL ensemble are found for subsequent evolutionary sampling (Table 1). Limited information, introduced by stochastic thresholding of test data, might lead to a performance threshold that is impossible to overcome theoretically. For the combined version of sampling and prior, a slightly worse result is found for $\mu_{D}$ and $\mu_{\rho}$, whereas the origin deviation is comparable to the ensemble mean and significantly better than the center of mass.
Even though parameter estimation does not improve with the network ensemble and subsequent evolutionary sampling, the rationale for training the ensemble is founded on the anticipation of achieving more stable predictions and providing variance estimation for the DL-Prior to optimally initialize the subsequent sampling. Further, the ill-posedness of the problem allows multiple parameter combinations to result in similar tumors. The following section examines this premise by comparing the resulting tumor concentrations.
@@ 3.2.3 Estimated Tumor Concentration
Subsequent sampling decreases the error to the ground truth of the estimated tumor concentration as shown in Table 2. This highlights the synergistic combination of sampling with a DL-Prior. Further, the combined version of CMA-ES sampling and DL-Prior achieves a significantly better Dice-score than the DL-Prior alone over the whole range of tumor concentrations (Figure 4). Within the clinically relevant range, up to 40% tumor concentration, a Dice overlap of 95% is found, further demonstrating the superiority of combining DL with traditional sampling.
@@ 3.2.4 Runtime comparison
To guarantee the integrity of our comparative analysis regarding time, we have included performance metrics for the specific subset of samples utilized in the temporal comparison, as detailed in Table 3. Sampling enhanced with a DL-Prior demonstrates improved results in terms of the Dice. However, the observed differences between these methods fall within the range of the method-specific error margins. These observations suggest that both approaches converge towards a similar minimum, legitimizing the time comparison of Figure 5.
Figure 5: (a) The mean likelihood over time for Naive Sampling and DL-Prior + Sampling is shown. Shaded areas represent standard error over different patients. DL-Prior + Sampling consistently yields higher likelihoods, with a significant difference around the clinically relevant 2-hour mark, resulting in a likelihood of 0.85 $\pm$ 0.03, while Naive Sampling only reached 0.41 $\pm$ 0.05. This translates to a p-value of $1.2\times 10^{-5}$ and an effect size of 2.0. (b) The mean convergence time of our DL-inference (orange) and our combination of classical sampling with DL-Prior (blue) is compared to sampling without prior (purple, dashed). The paired t-test results in an effect size of 1.0 and a p-value of 0.003
Table 3: Tumor concentration comparison between DL-Prior with subsequent sampling and sampling without a prior for a subset of patients. The Dice-score at different thresholds (0.01, 0.1, 0.5) is given, and no substantial difference was found. Simulation time differences are shown in Figure 5.
@@ 4 Validation on Patient Data
The quantitative ability of different methods to fit real tumor segmentations is compared in Table 4. The Dice-score is compared for CET and edema regions. It is clearly visible that sampling improves the simulation fit dramatically. The differences in mean Dice-score between naive sampling and prior included sampling are small, but those results do not allow for statistical statements. Therefore, we qualitatively analyze the results case-specific in Figure 7. Regarding the runtime, a clear advantage is seen for DL-Prior + sampling.
Figure 6: Real patient example with different inversion strategies. The first image shows the ground truth tumor segmentation before surgery. Light green shows the edema region, while dark green represents the CET, which is included in the following images as a black outline. The mean likelihood development over time comparing DL-Prior sampling to naive sampling is plotted. The margin shows the minimum and maximum values for each CMA-ES epoch, and the vertical lines depict the time points of convergence. The likelihood for the other methods is 0.39 for LMI, 0.23 for DL, and 0.13 for DL ensemble.
Table 4: For nine real patients, the Dice-score to the tumor segmentation is compared for different algorithms, thresholded at the $\overline{c^{th}}=\{0.25,0.675\}$. Additionally, the convergence time is reported for each method. It becomes apparent that naive sampling and sampling with a prior reach similar Dice-score with significant time differences in runtime.
@@
Naive Sampling 0.70 $\pm$ 0.03 0.70 $\pm$ 0.02 (16.3 $\pm$ 1.7) h*
DL 0.39 $\pm$ 0.06* 0.44 $\pm$ 0.04* $<$ 1 min*
DL Ensemble 0.44 $\pm$ 0.08* 0.47 $\pm$ 0.05* $<$ 1 min*
LMI 0.34 $\pm$ 0.09* 0.55 $\pm$ 0.04* $<$ 1 min*
Our method generally demonstrates strong performance, as illustrated in Figure 5, which also shows significant runtime improvements. However, there are occasional instances where it encounters challenges. We analyze various example scans in Figure 7 to assess the types of tumors that our method accurately characterizes and to identify specific scenarios where our approach falls short, highlighting potential for future research. For the small tumor (a), the likelihood plot indicates a strong advantage for including the DL-Prior convergence, which is reached much faster than with naive sampling. The naive sampling approaches the optimal prior region only after exploring a larger space, while the diffusion and growth rate stay close to the initial prior. For the larger tumor (b), the prior estimate of the origin is close to the final estimate of the naive process. The prior sampling early advantage diminishes over time, as the DL-Prior restricts the diffusion rate to decrease further, which seems beneficial for this tumor. However, for the example with large mass effect (c), all methods struggle to fit the tumor, which is expected, as the mass effect is not part of the physical tumor model. Due to the insufficient tumor model, the DL-Prior becomes obsolete, and no improvement in likelihood is achieved. Neither method fits the scans well in this case.
In summary, the introduction of a DL-Prior helps the sampling strategy to converge faster to the optimal solution (Figure 5, 7 a). In some cases, the underlying tumor model is clearly oversimplified (7 b), and in a few other cases, the prior can lead the sampling in the wrong direction (7 c). The last problem could be overcome by relaxing the prior constraint over time.
Figure 7: Evaluation of our method focusing on multiple scans of varying sizes and mass effects, particularly analyzing challenging cases. The final tumor concentration predictions are shown in the first row. Below, the time developments of $\mu_{D}$, $\mu_{\rho}$, the likelihood, and the tumor origin are shown. The DL and ensemble results are shown in red and orange. It becomes apparent that for small tumors without mass effect (a), the DL-Prior estimation provides a solid estimate leading to faster convergence. Large mass effects (c) can result in a failure of the network, constraining the sampling process into a suboptimal parameter set, which becomes apparent concerning the origin in this case.
@@ 5 Discussion
A new DL-based method is introduced, operating solely in patient image space compared to previous atlas-based methods [14] by explicitly incorporating patient anatomy into the network. With our method, a better parameter estimation is found (Table 1). Additionally, a better Dice overlap to the ground truth concentration is observed (Figure 4). We assume that the abundance of registration errors between the patient and the atlas space causes this improvement.
@@ 6 Outlook
Finally, an evaluation of large-scale clinical data is needed. As the availability of longitudinal data is highly limited due to the rapid surgery after detection for most patients, other validation metrics are considered. Prediction of tumor recurrence represents one of the most promising approaches. Besides the influence of treatment procedures like surgery and radiotherapy, it was shown that a certain degree of recurrence prediction is possible [43, 44, 45]. Showing the ability to predict recurrence with a physically constrained model would provide an excellent argument for proceeding with clinical trials testing model-based radiotherapy planning.
7 Conclusion
This study introduces a novel framework that synergistically combines deep learning with evolutionary sampling to enhance brain tumor modeling. By leveraging a DL ensemble for initial parameter estimation and incorporating it into a high-precision evolution strategy, we achieve a fivefold acceleration in convergence time and a Dice-score of 95%. This approach significantly constrains the sampling parameter space, offering a powerful and efficient solution for individualized radiotherapy planning with the potential to substantially improve patient outcomes.